Parallel Testing of LBC-Red and DX Cell during LUCIFER observing

17 February 2010 UT

Observer: JHill (Tucson)

Summary

JMH completed the first phase of the on-sky tests with LBC-Red and the DX Cell. A short summary of the results:

a) Both LBC-Red and the DX cell seem to be working OK.

b) I have updated the DX collimation table for LBC-Red. (Maybe it is now better than the one for LBC-Blue.)

c) I made the first tests that small repointings with the cell are repeatable. This was limited because I didn't have the Red Rotator running. The repeatability was already known to be OK from daytime tests.

d) You are able to use Red for LBC focussing tests tomorrow night (Feb 18 UT). Just remember that it has not been extensively tested, so there could be problems that we have not discovered.

e) I have not co-pointed Red and Blue, so that adjustment is needed if you want to observe with both at high elevation.

f) More work is scheduled in the DX cell for Thursday during the day. This probably makes it work even better, but has a small chance of making DX inoperable on Thursday night.

g) If you wish, Andrea (or even John) can change the hardwired flag so that Blue is the Guiding Master and Red is the Slave. The DX cell should now be free of the jitters that were making some Blue images elongated. As discussed in a previous email, the danger to switching the flag is that you may not always have a U-band guide star to guide the mount.

Details

Preparations

See NewHardpointCellCheckout for the broad test plan and other engineering tests on the DX hardpoints.

Thanks to John M. for remembering to remove the Red L1 cover in the afternoon.

I have changed redchannel.conf to have LBC-Red use the internal rotator trajectory and not the trajectory from TCS. This is so I can take LBC-Red images while Vincenzo et al. are observing with LUCIFER. And while I was there, I updated the TAI2UTC conversion from -33 to -34 sec.

OBs from /home/LBTO/Calib_OBs/

src/lbc/actions.c decides which side is the tracking slave. Previously it was always Red unless Red was disabled.

Startup

03:34 lbckill.sh/lbcstart.sh to be sure that the new Red configuration file was read.

03:40 Turn on LBC, Red and Blue. Blue Science failed to turn on. Did not connect to TCS.

03:42 RB_rVfastextra 034234 --- was with the Right LUCIFER collimation model which is empty.

03:50 JMH authorizes Right side as LBC, This stops the telescope tracking and screws up the LUCIFER spectroscopy observation. I believe this would have been OK in next month's TCS, but was not safe to do in Build 35.

Focussing

03:56 Now with the historical LBC-Red collimation model
RB_rVfastextra 035538 - seems like it is trailed?
RB_rVfastextra 040219 - EL=35 degrees
RB_rVfastextra 040405 - dZ = -0.7 mm (corrections sent by hand on PSFGUI)
RB_rVfastextra 040618 - peculiar illumination pattern
RB_rVfastextra 040959
send Z8 +1000
RB_rVfastextra 041314 - coma is worse
send Z8 -3000
RB_rVfastextra 041610 - coma is gone, but strange pupils
dZ = -0.1 mm and Z8 -500

The GUI is reporting R-BESSEL when I am expecting r-SLOAN. So I am suspicious that there is a filter wheel problem causing the peculiar pupil viginetting that I see. JMH: The following day it seems obvious that the vignetting has a vertical edge and must be a problem with the Red shutter. And it turns out that somebody changed the OB long ago to have R-BESSEL as the Red filter - They should have changed the OB name!

04:22 Turn Off LBC, then Turn On. Blue Science is OK this time.

RB_rVfastextra 042624 - still R-BESSEL but the pupils look OK now.
Z4 +2962 Z8 -843 from LBCFPIA applied manually
RB_rVfastextra 042930 at EL=38 deg
Z4 +792 Z6 406 Z8 -221 from LBCFPIA applied manually
RB_rVfastextra 043246 - well collimated pupils.

In focus

04:35 dZ = +0.8 mm to move to the focussed position.

04:36 10images5secV starting with 043624 with some manual collimates along the way

04:42 10images5secV 6x exp (30 sec) starting with 044254. 044651 has a LUCIFER dither in it.

Net collimation corrections: dX=-3.042 dY=0 dZ=-0.657 dRX=0 dRY=+32.68
I have updated the global offsets in DXPMLBCCollimation.dat accordingly.

The collimations were updated on 20100222 to reflect a new magic position.
Net collimation corrections: dX=-3.042 dY=-0.500 dZ=-0.157 dRX=0 dRY=+32.68

Focussing

RB_rVfastextra 050130 - but John wiped out his new globals adjusting dZ
RB_rVfastextra 050448 - has a dithering offset doubling the pupils
Z4 -989
RB_rVfastextra 050713 at EL=44 deg
Z4 +464 Z7 +361
RB_rVfastextra 050926
Z4 +677 dZ=+0.8 mm

In focus for pointing offset tests

05:12 10images5secV 12x exp (60 sec) starting with 051233
Guidestar is elongated and drifting. EL=45 deg.
The drift was later found to be because the rotator was not running.

Each image starts with the PSF Tip-Tilt Pointing Offsets at 0,0 arcsec.

051725 5 arcsec Tip at 30 sec, 5 arcsec Tilt at 50 sec - Offsets appear not exactly orthogonal.

051856 -5 arcsec Tip at 35 sec, -2 arcsec Tilt at 50 sec - Must have included a LUCIFER dither.

052028 -5 arcsec Tip at 30 sec, 0 arcsec Tip at 45 sec - There are only two images, so the offsets were precise at the arcsec level.

052200 +5 arcsec Tilt at 20 sec, 0 arcsec Tilt at 40 sec - Only two images. Spacing is 21 pixels = 4.7 arcsec.

Focussing

RB_rVfastextra 055438 at EL=52. Wavefront is still good from a while ago.

06:11 Put the red configuration file back to normal.

06:17 Turn off LBC. lbckill/lbcstart

-- JohnHill - 17 Feb 2010
Topic revision: r4 - 22 Feb 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback